You are considering investing in Nuran Security Services. You have been able to locate the following information on the firm: total assets are $18 million, dividend is 0.8 million, retention ratio is 60%, and debt-to-equity ratio is 1.5 times.

Calculate the ROE
Answer #1

Dividend = $ 0.8 million and Retention Ratio = 60 %

Dividend Payout Ratio = (1-Retention Ratio) = (1-0.6) = 0.4 or 40 % of Net Income

Therefore, Net Income = Dividends / Dividend Payout Ratio = 0.8 / 0.4 = $ 2 million

Total Assets = $ 18 million and Debt to Equity Ratio = 1.5

Therefore, Debt (1.5/2.5) x 18 = $ 10.8 million and Equity = (1/2.5) x 18 = $ 7.2 million

ROE = (Net Income / Equity) x 100 = (2/7.2) x 100 = 27.78 %